Number if people
108
44
206
85
57

Question 18 Which one of the following statements is incorrect? (1) The mean and the median are equal. (2) The mean is less than the median. (3) The mode is zero. (4) There is no mode. (5) None of the above.
Question 19 What is the sample standard deviation? (1) 64:25: (2) 57:46: (3) 4127:50: (4) 3302: (5) None of the above. 21
Question 20 The coefficient of variation is: (1) 64:25%. (2) 57:46%. (3) 4127:50%. (4) 3302%. (5) 100%.
1
Expert's answer
2020-04-06T17:47:36-0400

Question 18

Answer: (1) The mean and the median are equal;

(2) The mean is less than the median; (3) The mode is zero

are incorrect. This is because the mean =100 and median =85

For (1) and (2) The mean is not equal to the median neither is it less than the median. For statement (3) there is no mode in the dataset above.

Question 19

Sample standard deviation

"s=\sqrt \frac {\sum {(x-\bar x)^2 }} {{n-1}}"

="\sqrt {\frac {(44-100)^2+(57-100)^2+(85-100)^2 +(108-100)^2+(206-100)^2}{5-1}}=64.25"

Answer: (1) 64.25.

Question 20

Coefficient of Variation

="\frac {s} {\bar x} *100"%


="\frac {64.25}{100}*100\%=64.25\%"

Answer: 1) 64.25%.